OpenStudy (anonymous):

A cube is inscribed within the sphere such that all the vertices of the cube touch the sphere. The volume of the cube is equal to 2×963√, what is the radius of the sphere?

OpenStudy (anonymous):

you need calculas unless you know what the theta of the angle of the radius to the sectum is?

OpenStudy (anonymous):

do you know the angle?

OpenStudy (anonymous):

diameter of the sphere is the diagonal of the cube so the diameter = diagonal = sqrt(s*s + s*s + s*s) and then r=d/2